Grade 11Physical Chemistry

For the rchemical equation,N2+3H2 gives 2NH3 calculate the mass of NH3 formed if 2*10*3g N2 reacts with 10*3g of H2.will any of the two reactants remains unreacted?If yes,which one and would be its mass?

kavya sharma
9 Years agoGrade 11

1 Answer

Devender Lour
9 Years ago
N2+3H2=2NH3So 28g N2 will react with 6g H260g N2 will react with 6/28×60g H2 =12.8574gH2 left unreacted=30-12.8574=17.1426gNow we notice that 28g N2 gives 34g NH360g N2 will give=34/28*60=72.8574g NH3